OpenStudy (anonymous):

log7 + log (n - 2) = log 6n Can someone help? I really REALLY do not understand logarithms

OpenStudy (anonymous):

is this exactly like it's written in your book? There aren't any small numbers by the logs??

OpenStudy (anonymous):

I don't have a book. I am taking my class online.

OpenStudy (anonymous):

oh ok. But there are no small numbers?

OpenStudy (anonymous):

But yes. That is it exactly. In fact I copy/pasted.

OpenStudy (anonymous):

Ok then we are lucky! Do you know what it means to do 10^n? Doing a log is just doing that backwards

OpenStudy (anonymous):

Casey, not really. I'm super duper lost.

OpenStudy (anonymous):

ok ok no worries.

OpenStudy (anonymous):

Do you know what exponents do?

OpenStudy (anonymous):

Yeah. They denote how many times you multiply a number by itself. 5^3 would be 5*5*5

OpenStudy (anonymous):

perfect. So 10^3 is 10*10*10, right? and 10^n just means we multiply 10 together "n" times, and "n" could be any old number we end up putting in there

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So now logarithms... Just like the reverse of multiplying is dividing, the reverse of 10 to the power of something (10^n) is the log of that something (log n)

OpenStudy (anonymous):

so let's say we want to multiply 10 together a certain number of times to get 10,000... and we want to know how many times we have to multiply 10 together to get to 10,000

OpenStudy (anonymous):

can you tell how many times it is?

OpenStudy (anonymous):

Okay. 4 times.

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

So we right that like this

OpenStudy (anonymous):

\[10^x=10,000\]

OpenStudy (anonymous):

and because we know that the log is the reverse of 10^something, we'll do the log of both sides to get x by itself so we can see the answer

OpenStudy (anonymous):

\[\log(10^x)=\log(10,000)\]

OpenStudy (anonymous):

because the log and the 10^ are reverses, they cancel each other and go away

OpenStudy (anonymous):

\[x=\log(10,000)\]

OpenStudy (anonymous):

and you can do the log of 10,000 on your calculator and see that it's 4, like you already did

OpenStudy (anonymous):

Oh. :o

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Yes. :o So now, how do I add logs together?

OpenStudy (anonymous):

ok well... we don't have to! That's the great thing! because 10^something and log cancel out, you can do this for you original question:

OpenStudy (anonymous):

Now you lost me.

OpenStudy (anonymous):

\[10^{\log(7)}=10^{\log(n-2)}=10^{\log(6n)}\]

OpenStudy (anonymous):

just wait, you'll get it.

OpenStudy (anonymous):

So they all equal the same thing??

OpenStudy (anonymous):

just take 10^ each term... like I did. Then, the 10^ and the log cancel, right? So they go away

OpenStudy (anonymous):

and you get \[7+n-2=6n\]... where there are no logs to confuse us and we can just find n with normal old algebra

OpenStudy (anonymous):

Oh! :o

OpenStudy (anonymous):

make a little bit of sense?

OpenStudy (anonymous):

yeah. So n would be 5/7?

OpenStudy (anonymous):

make sure you subtract the n from both sides, so the right side should be 5n at that point, then divide by 5

OpenStudy (anonymous):

Oh its subtracting. Oops XD so n would equal 1.

OpenStudy (anonymous):

perfect. Do you get it? Or do you want to do one more similar problem?

OpenStudy (anonymous):

Can we do this one; log(5)16 – log(5)2t = log(5)2?? The numbers in the parenthesis are the little numbers

OpenStudy (anonymous):

ah ok I thought we were going to go there. Ok, so no big deal, it's just as easy as the last problem. We just need to know what those numbers mean. Every log has a base. Plain old log is actually \[\log_{10}\]but we use it so often we usually don't write the 10

OpenStudy (anonymous):

so remember how \[\log_{10}\]was 10^n backwards? Well, \[\log_{5}\] is just 5^n backwards

OpenStudy (anonymous):

Okay. So its always going the be b^n backwards? b for the base

OpenStudy (anonymous):

exactly... and we usually call the base b. You're way ahead of me :)

OpenStudy (anonymous):

So, knowing that, how could you solve this problem?

OpenStudy (anonymous):

5^16 - 5^(2t) = 5^2???

OpenStudy (anonymous):

you've got the right idea, but remember that the 5^ and the log(5) cancel. Try one more time?

OpenStudy (anonymous):

16-2t=2???

OpenStudy (anonymous):

there it is!

OpenStudy (anonymous):

So with logs I'm just taking their opposite, getting rid of the base and solving the exponents?

OpenStudy (anonymous):

yah, that's it exactly.

OpenStudy (anonymous):

Try this one... if you get it I'm sure you've got it under control

OpenStudy (anonymous):

\[\log_2(4x)=4\]

OpenStudy (anonymous):

hrm. log(2) is just 2^ so 2^(4x) = 4 but the log and the 2^ cancel so its just 4x = 4 so x = 1????? Did I do that right???

OpenStudy (anonymous):

@caseyrt

OpenStudy (anonymous):

just about. You did the left side exactly right. But remember, what you do to one side you have to do to the other side. so the right side needs to have 2^4, too. What do you get now?

OpenStudy (anonymous):

Oh so it should be 4x = 2^4? then 4x = 16 x=4???

OpenStudy (anonymous):

Winner winner! You nailed it!

OpenStudy (anonymous):

Okay thank you XDD You wouldn't mind checking my work on logs in the future would you ??? XDD

OpenStudy (anonymous):

no problem. I'm not on here super often, but you can message me on fb if you want. I'm just at /caseyrt

OpenStudy (anonymous):

Thank you c:

hero (hero):

log(a) + log(b) = log(ab) log7 + log (n - 2) = log 6n log(7) + log(n - 2) = log(7(n - 2)) log(7(n - 2)) = log(6n) log(7n - 14) = log(6n) 7n - 14 = 6n 7n - 6n = 14 (7 - 6)n = 14 n = 14

hero (hero):

@SavannahWillett and @caseyrt any objections?

OpenStudy (anonymous):

oh shoot.

hero (hero):

Remember to always use the "check method"

OpenStudy (anonymous):

wait, what do you mean?

OpenStudy (anonymous):

@Hero, what is "check method"?

hero (hero):

Check: n = 14 log(7) + log(14 - 2) = log(6(14))

OpenStudy (anonymous):

ok I see what you mean. I thought there might have been a button somewhere I was supposed to use to have someone check me. I feel so bad. I'll message her and fix it. Thank you

OpenStudy (anonymous):

That kind of makes sense @Hero